Web14 Outage (Resolved)
  • Priority - Critical
  • Affecting Server - web14
  • We are investigating why the to sever has gone down.

  • Start Time - 19/07/2019 18:23 - 19/07/2019 19:00
  • Last Updated - 22/07/2019 10:53
Server unavailable (Resolved)
  • Priority - Critical
  • Affecting Server - web06
  • Web06 has become unavailable. Techs are attending to the issue onsite currently. No ETR is available at this stage

    The server is now back online

  • Start Time - 22/07/2019 10:45 - 22/07/2019 10:52
  • Last Updated - 22/07/2019 10:53
Web11 outage (Resolved)
  • Priority - Critical
  • Affecting Server - web11
  • We are investigating reports of the server being down.

  • Start Time - 19/07/2019 16:11
  • Last Updated - 19/07/2019 18:23
Web14 Outage (Resolved)
  • Priority - Critical
  • Affecting Server - web14
  • Web14 is currently down and we are investigating the cause of the issue.

    [UPDAT 11:35AM WAST]

    We have managed to bring the server back online and we are continuing to monitor the issue.

  • Start Time - 03/07/2019 11:28 - 03/07/2019 11:39
  • Last Updated - 03/07/2019 11:40
Web09 outage (Resolved)
  • Priority - Critical
  • Affecting Server - web09
  • We have received reports of web09 being down. Our technicians are currently on site investigating the issue.


    [UPDATE 5:53]

    Our technicians managed to resolve the issue and we are monitoring the server closely.

  • Start Time - 06/06/2019 17:45 - 06/06/2019 17:53
  • Last Updated - 06/06/2019 17:53
Web14 Outage (Resolved)
  • Priority - Critical
  • Affecting Server - web14
  • We have reports of the server being offline. We are investigating the issue and have people on site working to resolve the issue.

    [UPDATE]

    We have got the server back online and we are monitoring the server.

  • Start Time - 29/05/2019 16:19 - 29/05/2019 16:33
  • Last Updated - 29/05/2019 16:34
WP03 Offline (Resolved)
  • Priority - Critical
  • Affecting Server - wp03
  • We are currently investigating the issue into why the server has gone offline. Our technicians are on site working to resolve the issue ASAP.

    You can be kept up to date via the server status page here - https://clients.netorigin.com.au/serverstatus.php

  • Start Time - 27/03/2019 16:30 - 27/03/2019 18:00
  • Last Updated - 07/04/2019 12:08
WP05 (Resolved)
  • Priority - Critical
  • Affecting Server - wp05
  • We have now resolved the issue and the server is back online. We have a team investigsting the cause of the Downtime.






    We are currently investigating the issue on the server. Updates will be provided as we find out more.

  • Start Time - 23/03/2019 14:45 - 23/03/2019 00:00
  • Last Updated - 23/03/2019 15:06
Web07 (Resolved)
  • Priority - Critical
  • Affecting Server - web07

  • 12:11; PM 30/01/2019
    We are currently aware of email issues on web07 and our system administrators are working to fix the issue.

  • Start Time - 30/01/2019 12:11
  • Last Updated - 31/01/2019 14:53
Datacenter Power Issue (web07 server issue) (Resolved)
  • Priority - Critical
  • Affecting System - Network
  • DC staff are currently investigating an issue affecting connectivity and power to servers within the facility. An update will be provided when further information is available.


    UPDATE 9:39 AM
    Most Services are now back online. The DC staff are investigating the cause of the outage.

    Issues are still being investigated with web07 and web09


    UPDATE 12:49PM
    web07 and web09 continue to experience hardware issues. System admins are working to restore services to both servers as soon as possible. We are working on implementing two new servers to restore backups to in the event that the servers cannot be brought back online.

    UPDATE 4:10PM
    web09 is now back online with a high load, however this is expected to stablise shortly. System admins have been unable to restore services to web07 at this point. A new server has been provisioned with system admins working to restore data to the new server from our backups. An ETA on expected completion time for the backup restoration will be provided in the next couple of hours.

    UPDATE 6:04PM
    A backup for web07 is currently being restored to a newly built server. There is a considerable amount of data to be restored, and the current estimated completion time for the backup is 24 hours.

    UPDATE 12:04AM
    The backup for web07 continues to be restored. At this point the estimate completion time remains the same.

    UPDATE 8:07AM
    The backup for web07 continues to be restored, however the restoration speed has slowed considerably overnight. At this point the backup restoration is expected to be completed by Sunday evening.

    UPDATE SUNDAY 9:54AM
    The backup for web07 continues to be restored. An updated ETA for the restoration to be completed is 11:30PM (AWST).

    UPDATE Monday 1:45AM
    The restore has completed and the server is back up and running. An investigation into the duration of the restore process will be under taken with the software developer and further investigation into the hardware fault is ongoing. 

  • Start Time - 25/01/2019 09:27 - 28/01/2019 01:40
  • Last Updated - 28/01/2019 01:45
Web15 RAM SWAP (Resolved)
  • Priority - Critical
  • Affecting Server - web15
  • The RAM has been added and the server should now be stable.

  • Start Time - 27/12/2018 10:29 - 27/12/2018 10:41
  • Last Updated - 27/12/2018 10:41
Web15 (Resolved)
  • Priority - Critical
  • Affecting Server - web15
  • The load have come back down and we are monitoring the server.

  • Start Time - 27/12/2018 09:00 - 27/12/2018 09:19
  • Last Updated - 27/12/2018 09:19
Drive Consistency Check - Server Maintenance (Resolved)
  • Priority - Critical
  • Affecting Server - web11
  • The Web11 server will be taken offline for a period within this window to perform a consistency check over the drive array after a statistical error was detected. The server will be unresponsive during the period of this check and will automatically resume normal operation once completed.

     

    Maintenance has now completed. 

  • Start Time - 27/12/2018 03:30 - 27/12/2018 06:00
  • Last Updated - 27/12/2018 05:21
Web15 Unresponsive (Resolved)
  • Priority - Critical
  • Affecting Server - web15
  • The server load is currently high again and we are investigating the cause of this.

  • Start Time - 17/12/2018 10:17
  • Last Updated - 18/12/2018 09:26
Web07 Unresponsive (Resolved)
  • Priority - Critical
  • Affecting Server - web07
  • The server web07 is currently unresponsive. We have a system admin investigating and an update will be provided in due course.

    A drive issue was detected and a consistency check of the array completed to ensure no damage to data occured. The server is now back in normal operation. 

  • Start Time - 12/12/2018 12:30 - 14/12/2018 13:31
  • Last Updated - 14/12/2018 05:35
Network Connectivity from some ISPs (Resolved)
  • Priority - Critical
  • Affecting Other - Network
  • We are recieving reports from multiple users of issues connecting to anything within our network. This has been reported to our upstream provider who appear to be experiencing an issue with a fibre cut interstate. At this point we do not have an ETR.

  • Start Time - 15/10/2018 12:42 - 15/10/2018 14:00
  • Last Updated - 15/10/2018 18:57
WEB14 (Resolved)
  • Priority - Critical
  • Affecting Server - web14
  • We currently have an outage on this server. We have a team investigating the issue.

    UPDATE: The server has been rebooted and we are monitoring the server closely.

  • Start Time - 30/07/2018 16:23 - 31/07/2018 00:00
  • Last Updated - 03/10/2018 08:54
WEB13 (Resolved)
  • Priority - Critical
  • Affecting Server - web13
  • Currently, we are investigating reports of web13 not resolving.

    We have a team on site working to resolve the issue.

     

  • Start Time - 28/07/2018 21:43 - 28/07/2018 21:55
  • Last Updated - 28/07/2018 21:58
Intermittent Service Issue - Managed VDS serv (Resolved)
  • Priority - Critical
  • Affecting Other - Managed VDS services
  • An issue with intermittent response effecting reliability and performance on Managed VDS clients is currently being investigated by an upstream provider.

    The upstream provider has identified and rectified the issue effecting the performance for Managed VDS clients. A full analysis is being undertaken to understand and rectify the cause

  • Start Time - 02/07/2018 11:30 - 02/07/2018 14:41
  • Last Updated - 02/07/2018 12:43
.au registry maintenance (Resolved)
  • Priority - Critical
  • Affecting Other - .au Domains
  • Maintenance of the .au registry
    There will be scheduled downtime of the .au registry between 8am Saturday 30th June to 8pm Sunday 1st July, AEST.

    During this period .au domain name registrations, renewals and transfers will not be available. Additionally change of Registrant information or DNS delegation will not be able to be processed for a minimum of 48 hours.

  • Start Time - 30/06/2018 06:00 - 01/07/2018 22:59
  • Last Updated - 02/07/2018 08:55
Web07 High CPU (Resolved)
  • Priority - Critical
  • Affecting Server - web07
  • We currently have a team investigating the issue were web07 has a constantly high CPU usage.


  • Start Time - 24/05/2018 13:51
  • Last Updated - 24/05/2018 14:44
Web06 Outage (Resolved)
  • Priority - Critical
  • Affecting Server - web06
  • Our team are currently investigating an issue effecting the server web06 being unresponsive.

    Update: The system is currently undergoing a disk consistency check due

    Update (712am): The consistency check has completed and the server is running. We are monitoring further to see if this has resolves the intermittent responsive issues that were being experienced prior

  • Start Time - 08/05/2018 06:30 - 08/05/2018 07:20
  • Last Updated - 08/05/2018 21:28
Disk Maintenance (Resolved)
  • Priority - Critical
  • Affecting Server - web11
  • The web11 server is currently undergoing maintenance on it's disk array and is offline as a result.

    Update: The maintenance has completed and the server is back up and responding to requests. 

  • Start Time - 05/05/2018 14:30 - 05/05/2018 17:00
  • Last Updated - 05/05/2018 16:00
Server Unresponsive (Resolved)
  • Priority - Critical
  • Affecting Server - mail01
  • The server mail01 is experienced an unplanned outage and is currently being investigated by our systems team.

    Update - The server has been rebooted and is back operational. Investigations are continuing into the cause. 

  • Start Time - 30/04/2018 12:46 - 30/04/2018 15:03
  • Last Updated - 30/04/2018 13:04
Intermittent SMTP connection issues (Resolved)
  • Priority - Critical
  • Affecting Server - mail01
  • We have received reports of intermittent SMTP connection issues to the server mail01 

    Some users have reported certian device types can connect fine while others report they're unable to connect using a variety of settings. This issue is currently being
    investigated further by our systems team to isolate the issue.

     

    Update: After further diagnosis, this issue appears to effect people using insecure SSL cyper methods after a software update. Support was re-enabled for older cyphers, but users are recommended to update their Operating System and Email Programs for security reasons. 

  • Start Time - 16/04/2018 08:55 - 16/04/2018 12:47
  • Last Updated - 16/04/2018 12:50
Server Update (Resolved)
  • Priority - Medium
  • Affecting Other - Multiple Servers
  • In order to apply critical security across a range of servers, a series of rolling reboots will be undertaken. Servers will be offline during this process for between 5 and 15 minutes as the updates are applied.

  • Start Time - 09/04/2018 09:10 - 16/04/2018 10:57
  • Last Updated - 09/04/2018 09:31
Pegasus (Resolved)
  • Priority - Critical
  • Affecting Server - SolusVM
  • We are currently experiencing issues with the host node server, and have a system admin on site investing the issue.

  • Start Time - 11/03/2018 11:49 - 11/03/2018 13:11
  • Last Updated - 11/03/2018 15:04
DDOS Mitigation (Resolved)
  • Priority - Critical
  • Affecting Server - web14
  • We have received advise from an upstream provider regarding a DDOS attempt to the Web14 server which is being mitigated. 

    During DDOS mitigation some legitimate attempts to access site and services may be interupted. 

    We will provide updates on the process as they become available from the upstream provider. 

     

  • Start Time - 08/02/2018 09:00 - 08/03/2018 10:00
  • Last Updated - 01/03/2018 08:55
Upstream DDOS (Resolved)
  • Priority - Critical
  • Affecting System - Network
  • Our upstream carrier has reported they are currently investigating a DDOS which is impacting services to downstream clients including ourselves. 

    Some services will be effected for the duration of this incident such as slow or no page loading. 

     

    Update: Scrubbing has now been implemented to isolate the traffic and stabalise the upstream network. We are seeing network usage return to normal. 

  • Start Time - 18/12/2017 19:45 - 18/12/2017 19:58
  • Last Updated - 18/12/2017 20:01
SSL Orders and Renewals (Resolved)
  • Priority - Medium
  • Affecting Other - SSL
  • Unfortunately, we are currently experiencing issues with new SSL orders and renewals. At this point in time, we cannot provide an ETA and our SSL issuer is currently investigating the issue.

  • Start Time - 04/12/2017 11:30 - 00/00/0000 00:00
  • Last Updated - 05/12/2017 08:13
Network Issue (Resolved)
  • Priority - Critical
  • Affecting Server - SolusVM
  • We are currently investigating a network issue. Initial reports indicate a fibre cut in Malaga which may impact some customer services.

    We will provide an updates shortly.

    Update - 12:28AM (AWST)
    A network admin is currently implementing a work around. We expect affected services to return to normal within the next half an hour.

     

  • Start Time - 27/11/2017 23:43 - 28/11/2017 20:00
  • Last Updated - 04/12/2017 12:10
Server Maintenance (Resolved)
  • Priority - High
  • Affecting Server - wp04
  • We will be performing hardware maintenance on the Managed WordPress server wp04.netorigin.com.au. The maintenance will result in approximately half an hour of downtime.


  • Start Time - 22/11/2017 18:00 - 22/11/2017 18:35
  • Last Updated - 22/11/2017 15:45
Server Unresponsive (Resolved)
  • Priority - Critical
  • Affecting Server - wp05
  • We are currently experiencing connectivity issues with the server wp05.myhost.net.au. We have a system admin investigating.

  • Start Time - 09/11/2017 15:49
  • Last Updated - 09/11/2017 16:14
web06 HTTP server issue (Resolved)
  • Priority - High
  • Affecting Server - web06
  • We are currently experiencing issues with Litespeed Web Server on web06.netorigin.com.au. 

    A system administrator is investigating.


  • Start Time - 06/09/2017 09:19 - 06/09/2017 09:39
  • Last Updated - 06/09/2017 09:21
Upstream Transit Connectivity (Resolved)
  • Priority - Critical
  • Affecting Other - Network
  • An upstream connection issue has been detected resulting in some intermittent connection issues from connections coming in over a specific transit provider. 
    This issue has been reported to the carrier for further investigation. 

     

    [Update 11:11am WST] The carrier has confirmed an issue on their network. Changes have been made to isolate this carrier from being used until the issue is addressed and stability returns. 

    [Update 11:30am WST] The carrier has identified the issue as a core router failure and are currently working to replace and restore the service to normal levels. 

    [Update 11:40am WST] The carrier has advised that this issue has now been resolved and will be returned to normal usage. 

  • Start Time - 18/08/2017 11:05 - 18/08/2017 11:40
  • Last Updated - 18/08/2017 22:07